OpenStudy (anonymous):

How is this inequality factored? 5/x-8 > 7/6x-1 (5/x-8) - (7/6x-1) > 0

OpenStudy (asnaseer):

you should first get rid of the denominators in the fractions by first multiplying both sides by x to get:\[5-8x>\frac{7}{6}-x\]then multiply both sides by 6 to get:\[30-48x>7-6x\]can you solve it from there?

OpenStudy (anonymous):

-42x+23>0 Then I find the zeros, correct?

OpenStudy (asnaseer):

your simplification is correct, but there is no "finding of zeros" here. you just need to rearrange the equation to get an inequality for 'x'.

OpenStudy (anonymous):

Oh ok, so the zeros are 8 and 1/6

OpenStudy (anonymous):

?

OpenStudy (asnaseer):

\[-42x+23>0\]so:\[23>42x\]rearrange as:\[42x<23\]and then:\[x<23/42\]

OpenStudy (asnaseer):

this is assuming your original equation was:\[\frac{5}{x}-8>\frac{7}{6x}-1\]

OpenStudy (anonymous):

oh, no the original equation is not that....i need to type it correctly.

OpenStudy (asnaseer):

or was it:\[\frac{5}{x}-8>\frac{7x}{6}-1\]

OpenStudy (anonymous):

5 7 - > - x-8 6x-1

OpenStudy (anonymous):

And i have to "Solve the inequality"

OpenStudy (asnaseer):

ok, so it is:\[\frac{5}{x-8}>\frac{7}{6x-1}\]

OpenStudy (anonymous):

Yes, I could not find the division bar

OpenStudy (asnaseer):

first thing to do would be to multiply both sides by (x-8)(6x-1) to get:\[5(6x-1)>7(x-8)\]then expand to get:\[30x-5>7x-56\]then move all x's to the left-hand-side and all constants to the right-hand-side:\[30x-7x>5-56\]\[23x>-51\]\[x>\frac{51}{23}\]\[x>2\frac{5}{23}\]

OpenStudy (asnaseer):

sorry I dropped the minus sign in the last two steps

OpenStudy (asnaseer):

it should end up as:\[x>-2\frac{5}{23}\]

OpenStudy (asnaseer):

I hope it all made sense to you.

OpenStudy (anonymous):

Ok so is the correct solution \[(-\infty,-51\div23)\]

OpenStudy (asnaseer):

\[(-\frac{51}{23}, \infty)\]

OpenStudy (asnaseer):

remember, if z is a whole number and z > -3, it means it cannot be -3, -4, -5, ..., etc. it must be -2, -1, 0, 1, 2, ...

OpenStudy (asnaseer):

greater than a negative number means it must be bigger than that negative number - which means it must move towards the right on the real number line.

OpenStudy (anonymous):

ok, i started over and i just can not get your answer

OpenStudy (asnaseer):